### 17 Mar 2026 v1

# SegviGen: Repurposing 3D Generative Model for Part Segmentation

Abstract
We introduce SegviGen, a framework that repurposes native 3D generative models for 3D part segmentation. Existing pipelines either lift strong 2D priors into 3D via distillation or multi-view mask aggregation, often suffering from cross-view inconsistency and blurred boundaries, or explore native 3D discriminative segmentation, which typically requires large-scale annotated 3D data and substantial training resources. In contrast, SegviGen leverages the structured priors encoded in pretrained 3D generative model to induce segmentation through distinctive part colorization, establishing a novel and efficient framework for part segmentation. Specifically, SegviGen encodes a 3D asset and predicts part-indicative colors on active voxels of a geometry-aligned reconstruction. It supports interactive part segmentation, full segmentation, and full segmentation with 2D guidance in a unified framework. Extensive experiments show that SegviGen improves over the prior state of the art by 40% on interactive part segmentation and by 15% on full segmentation, while using only 0.32% of the labeled training data. It demonstrates that pretrained 3D generative priors transfer effectively to 3D part segmentation, enabling strong performance with limited supervision.
Figure 1: SegviGen enables diverse and accurate 3D part segmentation by leveraging priors from large-scale 3D generative models. With substantially less training data, it produces high-fidelity segmentation results with sharp part boundaries and strong generalization across object categories.
footnotetext: Equal contribution   Project lead   ✉ Corresponding author
1 Introduction
Part segmentation provides explicit part-level structures of 3D assets, serving as a core primitive for 3D content creation pipelines and offering fundamental 3D perception capabilities for spatial intelligence. It enables a wide range of downstream applications, including part-level editing, animation rigging, and industrial uses such as 3D printing. However, existing methods often fall short in segmentation quality, producing erroneous regions and imprecise boundaries that limit their practical usability.
To this end, one line of work attempts to transfer the comprehensive 2D segmentation priors to 3D via 2D-to-3D lifting. Methods such as SAMPart3D [79] optimize 3D segmentation via 2D-to-3D distillation, but incur substantial computational and time overhead, and often yield blurry boundaries. In parallel, another set of methods [21, 86, 10] applies SAM [26, 52, 1] to obtain 2D masks of multi-view projected images, which are then back-projected and fused into 3D masks. However, these multi-view pipelines incur substantial runtime overhead, are sensitive to view coverage, and the back-projection and fusion step often introduces cross-view inconsistencies and imprecise boundaries.
Recently, another line of work [43, 88] moves toward native 3D part segmentation so as to remedy the inherent shortcomings of the aforementioned methods that leverage 2D segmentation priors. These methods predict segmentation parts directly in the native 3D space, explicitly enforce semantic and structural consistency, and are more efficient at inference. However, it is a typical requirement to collect large-scale training datasets with curated 3D part annotations, where fine-grained annotations are costly and inconsistent across sources in granularity, hierarchy, and boundary definitions.
In summary, the first line of methods suffers from a mismatch between 2D priors and 3D structure, while the second relies on costly training from scratch. Therefore, a more promising approach is to leverage a prior model that encodes both 3D structure and texture to perform segmentation. In particular, 3D generative models trained on large-scale unannotated 3D textured assets internalize rich part-level structure and texture patterns, providing a strong 3D prior over geometry and appearance. Such priors encourage part segmentation with sharper boundaries, while reducing reliance on dense part annotations and extensive task-specific training. This motivates us to ask: How can 3D generative priors be effectively transferred to part-level 3D segmentation to improve quality and data efficiency?
Motivated by this perspective, we propose SegviGen, a generative framework for 3D part segmentation that leverages the rich 3D structural and textural knowledge encoded in large-scale 3D generative models. Specifically, we formulate part segmentation as a colorization task that fully exploits the capacity of 3D generative models. SegviGen encodes the input 3D asset into a latent representation and uses it, together with the task embedding and query points, to condition the denoising process. The model is trained to predict part-indicative colors, along with reconstructing the underlying geometry. This formulation naturally accommodates additional conditioning signals, enabling SegviGen to flexibly support interactive part segmentation, full segmentation, and 2D segmentation map–guided full segmentation under a unified architecture. Notably, while the first two tasks are common settings, 2D segmentation map–guided full segmentation is uniquely enabled by SegviGen, supporting arbitrary part granularity and more precise segmentation that is critical for industrial applications.
Qualitative and quantitative results show that SegviGen consistently surpasses the prior state of the art, P3-SAM [43], while using only 0.32% of the labeled training data. On interactive part segmentation, it achieves the best performance across all metrics on PartObjaverse-Tiny [78] and PartNeXT [60], with a 40% gain in IoU@1, an important metric that reflects the model’s single-click accuracy. On full segmentation without guidance, SegviGen outperforms the best baseline by 15% in overall IoU, averaged across datasets. Our main contributions are summarized as follows:
  • We propose SegviGen, a unified multi-task framework for 3D part segmentation that effectively exploits the structural and textural priors encoded in pretrained 3D generative models, enabling accurate and efficient segmentation.
  • We reformulate 3D segmentation as part-wise colorization, where SegviGen predicts the colors of actiave voxel as part labels in a single generative process.
  • Extensive experiments show that SegviGen outperforms the prior state of the art by 40% on interactive part segmentation and 15% on full segmentation, using only 0.32% of the labeled training data, highlighting the effectiveness of transferring 3D generative priors to part segmentation.
2 Related Work
2.1 3D Part Segmentation
Traditional 3D part segmentation is typically cast as supervised semantic labeling on points or faces, using fixed part taxonomies provided by curated 3D segmentation datasets [46, 4, 7, 49]. Concretely, these methods [49, 70, 69, 71, 16, 34] typically combine a 3D feature encoder with a segmentation head to predict dataset-specific part IDs. However, the closed-world nature of both the label space and the training data limits generalization, making it difficult to transfer to unseen object categories or arbitrary, non-canonical part decompositions.
To alleviate this generalization bottleneck, recent works exploit 2D foundation models as transferable priors [51, 28, 26, 52, 2, 47] for 3D part segmentation. A common strategy adopts a render-and-lift pipeline: it segments multi-view renderings with promptable 2D models and then projects and fuses the masks back onto the 3D surface [55, 80, 76, 87, 77]. Despite being straightforward, this pipeline is often limited by incomplete view coverage and cross-view inconsistencies, which can lead to imprecise or blurred part boundaries after 2D-to-3D aggregation. Another line leverages distillation or feature projection to supervise 3D predictors with transferred 2D representations or pseudo-labels [58, 15]; however, it still inherits the 2D–3D domain gap and multi-view alignment issues, and typically entails longer optimization and training cycles.
Recognizing the scalability and reliability issues of 2D-to-3D lifting, recent studies have shifted toward native feed-forward 3D segmentation that predicts masks directly on 3D representations at inference time. Representative efforts for open-world part segmentation include training queryable 3D predictors with automatically curated supervision [44], learning continuous part-aware 3D feature fields for direct decomposition [38], and prompt-guided 3D mask prediction models [86], with more recent large-scale native 3D part segmentation models such as P3-SAM [43] and PartSAM [88] further scaling training on millions of shape–part pairs. Despite encouraging progress, these native 3D approaches are fundamentally bottlenecked by the availability of large-scale, high-quality 3D part annotations, and the inconsistency of part taxonomies and granularity across datasets often introduces supervision mismatch, ultimately weakening cross-domain generalization.
Figure 2: Pipeline of SegviGen. We reformulate 3D part segmentation as a conditional colorization task. During training, given a 3D mesh and its part-color ground truth, we encode both with a pretrained 3D VAE, add noise to the ground-truth latent, and then concatenate the geometry latent, noisy color latent, and point-condition tokens to form the final latent input. Conditioned on the sampled timestep and a task embedding, the multi-task flow transformer predicts the noise residual for flow-matching training.
2.2 3D Generative Model
The rapid progress of diffusion-based generative models [19, 54], together with the emergence of large-scale, high-quality 3D data collections [9, 8], has catalyzed a wave of 3D generative methods [39, 40, 41, 20, 56, 24, 82, 65, 29, 64, 75, 59, 62, 37, 67, 85, 53, 72, 45, 36, 11, 3, 5, 61, 17, 18, 14, 83, 63, 31, 81]. A prevalent route builds 3D assets through a 2D-to-3D pipeline: models first synthesize multi-view imagery and subsequently reconstruct the 3D geometry and appearance from these views  [40, 41, 56, 64, 75, 62, 59, 23, 50, 22], yet view-to-view discrepancies in the synthesized images can propagate and degrade the final 3D quality.
In contrast, a growing family of native 3D generative models learns directly in 3D latent spaces, typically pairing a variational autoencoder [25] with a diffusion transformer (DiT) [48] to perform denoising over compact latents  [82, 29, 67, 85, 33, 6, 13, 84, 57, 35, 66, 68, 32, 74, 30, 73]. By learning to generate in a compact yet expressive 3D latent space, these models encode rich structural and texture knowledge across large-scale 3D assets, providing a strong transferable prior for downstream 3D part segmentation. In particular, TRELLIS2 [73] introduces a field-free structured latent via an omni-voxel sparse voxel representation (O-Voxel) that jointly models geometry and appearance, enabling efficient generation with sharp, high-frequency textures that better preserve fine-grained part boundaries for 3D segmentation.
3 Methodology
We propose SegviGen, a unified multi-task framework for 3D part segmentation that supports three practical settings: interactive part-segmentation, full segmentation, and full segmentation with 2D guidance. To leverage the prior knowledge encoded in a pretrained 3D generative model, we cast 3D segmentation as a colorization problem. Specifically, we encode the input 3D asset into a compact latent that conditions generation, and optionally augment it with user interactions or a 2D segmentation map. Conditioned on these inputs, the model reconstructs the 3D asset while predicting colors for active voxels in the structured 3D representation, where each color corresponds to an individual part, yielding the final segmentation. Below, we begin by describing the underlying 3D generative model (Sec. 3.1), followed by our task reformulation (Sec. 3.2), and then detail the overall pipeline (Sec. 3.3).
3.1 Preliminary: Structured 3D Generative Model
Recent work [73] organizes each textured 3D asset into a sparse set of active voxels on a regular grid, where every active voxel stores geometry and texture features aligned in 3D. This formulation leverages a flexible dual-grid construction to robustly handle arbitrary topology while encoding physically-based material attributes jointly with geometry for faithful appearance modeling. Given the sparse omni-voxel representation, a Sparse Compression VAE (SC-VAE) maps each voxelized asset feature tensor $\mathbf{x}$ to a compact structured latent $\mathbf{z}_{1}=E_{\phi}(\mathbf{x})$ and reconstructs it via $\hat{\mathbf{x}}=D_{\theta}(\mathbf{z}_{1})$, yielding an expressive yet highly compressed 3D latent space. On top of these latents, a conditional flow-matching generator learns a time-dependent vector field $\mathbf{v}_{\psi}(\mathbf{z}_{t},t,\mathbf{c})$ under conditioning $\mathbf{c}$ by matching the constant velocity along linear interpolants:
$\mathbf{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\quad t\sim\mathcal{U}(0,1),\quad\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\mathbf{z}_{1},$ (1)
$\mathcal{L}_{\mathrm{cfm}}=\mathbb{E}\Big[\big\|\mathbf{v}_{\psi}(\mathbf{z}_{t},t,\mathbf{c})-(\mathbf{z}_{1}-\mathbf{z}_{0})\big\|_{2}^{2}\Big].$ (2)
This latent generative pipeline enables efficient synthesis of geometry- and texture-consistent 3D assets, and the resulting structured latents capture rich joint statistics of shape and appearance, providing a strong transferable prior for fine-grained 3D part segmentation.
3.2 Task Reformulation and I/O Representation
We reformulate 3D part segmentation as a color prediction problem in a structured 3D representation. This choice matches our base 3D generative model [73], which jointly parameterizes geometry together with appearance attributes such as color, material properties, and roughness in a unified representation. To maximize reuse of the pretrained generative prior, we avoid introducing an additional segmentation-specific attribute channel, which would increase modeling and optimization complexity, and instead express segmentation targets directly in color space, the most visually intuitive attribute. We consider three task settings with consistent input and output formats.
Interactive part-segmentation is formulated as binary part extraction: given 3D points indicating a target part, we supervise the model to color the selected part in white and the remaining regions in black. Full segmentation targets multi-part decomposition: we assign each part a distinct color from a randomly sampled color palette and supervise voxel colors accordingly. Importantly, correctness is defined up to a permutation of colors within each object; any one-to-one assignment between predicted colors and parts is considered valid. To reduce sensitivity to particular color choices, we use K=10K{=}10 independently sampled palettes per shape, providing multiple colorizations for the same underlying partition. Full segmentation with 2D guidance additionally conditions the model on a rendered 2D segmentation map: we first colorize the 3D parts and render the corresponding 2D segmentation map, and we then train the model to generate 3D voxel colors that are consistent with the color assignments in the 2D guidance. Overall, this formulation preserves a unified model interface across settings, enabling a consistent architecture and training pipeline.
3.3 Unified Multi-Task 3D Part Segmentation
Overall framework.
To fully leverage pretrained 3D generative models, we cast 3D part segmentation as a conditional part-wise colorization task in 3D latent space. Given an input asset $X$, a pretrained 3D VAE encoder $E(\cdot)$ produces a encoded latent $z=E(X)$, which helps specify the active voxel support and anchors generation to the underlying shape. For each task, we construct a part-wise colorized target and encode it into the same latent space to obtain $y$, following the task-specific scheme in Sec. 3.2. We then sample $\epsilon\sim\mathcal{N}(0,I)$ and $t\sim\mathcal{U}(0,1)$ to form a noisy interpolation
$y_{t}\;=\;(1-t)\,y\;+\;t\,\epsilon.$ (3)
A pretrained DiT-based backbone is fine-tuned to predict the noise residual conditioned on the noisy input $y_{t}$, the geometry latent $z$, the task condition $C$, and a learned task embedding $e_{\tau}$:
$\hat{v}_{\theta}\;=\;f_{\theta}\!\left(y_{t},\,z,\,C,\,e_{\tau},\,t\right).$ (4)
Training follows the conditional flow-matching objective
$\mathcal{L}(\theta)\;=\;\mathbb{E}_{X,\tau,t,\epsilon}\Big[w(t)\,\big\|\hat{v}_{\theta}\;-\;(\epsilon-y)\big\|_{2}^{2}\Big],$ (5)
where $w(t)$ is an optional timestep weighting.
Figure 3: Interactive part-segmentation results. We compare SegviGen with existing representative baselines, including Point-SAM [86] and P3-SAM [43]. In the figure, yellow points denote user clicks, while the predicted target part is highlighted in red. Leveraging priors from pretrained 3D generative models, SegviGen achieves more accurate results with sharper boundaries than prior methods, while requiring substantially less training data.
Condition Injection.
We adopt task-specific conditioning designs while maintaining a unified interface across settings. For interactive segmentation, user clicks in the UI provide an efficient and intuitive form of guidance. In our framework, each click is encoded as a sparse point token comprising its 3D coordinates and an associated feature vector. Since the 3D coordinates are already effectively encoded by RoPE within the attention layers, we omit the additional learnable input-level positional embedding used in prior designs [43]. Instead, all points share the same learnable feature vector $Q$ , which serves as the point token during both training and inference. Given point coordinates $\mathbf{u}=\{u_{i}\}_{i=1}^{m}$ with $u_{i}\in\mathbb{R}^{3}$, we form point-condition tokens
$Q\;=\;\big[\;\mathbf{q}(u_{1}),\ldots,\mathbf{q}(u_{m})\;\big],\quad\mathbf{q}(u_{i})\;=\;\big[\,u_{i}\,;\,\mathbf{e}_{p}\,\big],$ (6)
where $\mathbf{e}_{p}$ is a shared learnable feature appended to every point token. Conditioned on $Q$, the denoising model is instantiated as
$\hat{v}_{\theta}\;=\;f_{\theta}\!\left(y_{t},\,z,\,Q,\,e_{\tau},\,t\right).$ (7)
When the number of points is fewer than $10$, we pad the point tokens to a length of $10$ using zero coordinates and zero features. To preserve a single unified model, we keep this interface for full segmentation and 2D-guided full segmentation by providing $10$ padded tokens with all-zero coordinates and features.
For full segmentation with 2D guidance, we additionally provide a user-specified 2D segmentation colorization as guidance, allowing direct control over the desired part granularity and label palette. The guidance image is encoded into a sequence of conditioning tokens injected via cross-attention:
$p\;=\;g_{\phi}(I_{\text{guide}}),$ (8)
where $g_{\phi}(\cdot)$ denotes an image encoder. In this setting, denoising is conditioned on both the padded point-token interface $Q_{0}$ and the image guidance tokens $p$:
$\hat{v}_{\theta}\;=\;f_{\theta}\!\left(y_{t},\,z,\,(Q_{0},\,p),\,e_{\tau},\,t\right).$ (9)
Task Embedding.
To improve multi-task generalization within a single model, task identity is encoded as a continuous embedding and injected alongside the timestep signal. Let $\tau\in\{1,\dots,T\}$ denote the task index. A sinusoidal encoding is first computed from $\tau$,
$s_{\tau}\;=\;\mathrm{PE}(\tau)\in\mathbb{R}^{d_{f}},$ (10)
where $\mathrm{PE}(\cdot)$ follows the standard sinusoidal scheme. A lightweight MLP then maps $s_{\tau}$ to the task embedding
$e_{\tau}\;=\;\mathrm{MLP}_{\psi}(s_{\tau})\in\mathbb{R}^{d}.$ (11)
In parallel, the timestep $t$ is embedded as $e_{t}\in\mathbb{R}^{d}$. The final modulation vector used by DiT backbone is obtained by additive fusion,
$m\;=\;e_{t}\;+\;e_{\tau},\qquad\hat{v}_{\theta}\;=\;f_{\theta}\!\left(y_{t},\,z,\,C,\,m\right),$ (12)
where $m$ conditions the adaptive layers to jointly encode diffusion progress and task semantics. During training, samples from different tasks are interleaved and supervised with their corresponding $\tau$, encouraging the shared backbone to learn task-discriminative behaviors while preserving a unified parameterization.
4 Experiments
4.1 Setting
Table 1: Comparison of interactive part segmentation performance. We report IoU at different numbers of clicks, compared with Point-SAM [86] and P3-SAM [43] on PartObjaverse-Tiny [78] and PartNeXT [60].
Method PartObjaverse-Tiny PartNeXT
IoU@1 IoU@3 IoU@5 IoU@7 IoU@10 IoU@1 IoU@3 IoU@5 IoU@7 IoU@10
Point-SAM [86] 24.87 48.99 59.67 64.33 67.99 23.90 47.50 56.71 61.23 65.04
P3-SAM [43] 33.04 50.57 53.78 54.74 55.51 35.61 51.26 52.03 52.61 53.81
SegviGen 42.49 61.14 67.53 71.50 75.02 54.86 71.15 78.11 79.96 82.73
Figure 4: Full segmentation results. We compare SegviGen against a broad set of prior methods, where different colors indicate different segmented parts. From the results, SegviGen achieves high-accuracy full segmentation with sharp part boundaries using only 3D input. When 2D guidance is provided, it further allows explicit control over granularity and labels, enabling controllable, ultra-fine-grained 3D part parsing.
Implementation Details. We adopt Trellis.2 [73] as our base model, which is a 3D generative framework with a native and compact structured latent representation. For all experiments, the Tex-SLAT flow model is trainable, while the remaining SC-VAE is kept frozen. We adopt the AdamW optimizer [42] with a learning rate of $1\times 10^{-4}$. All experiments are conducted on 8 NVIDIA A800 GPUs, and the model is trained for 8 hours. Unless otherwise specified, the segmentation results shown in this paper are produced with 12-step inference.
Datasets. For training, we use the PartVerse dataset [12], which contains 12k objects with a total of approximately 91k annotated parts. For evaluation, we use PartObjaverse-Tiny [79], which contains 200 textured mesh objects, and a 300-object textured-mesh subset of PartNeXT [60].
Baselines. We compared our model’s performance on full segmentation between P3-SAM [43], Find3D [44], SAMPart3D [79], Partfield [38]. P3-SAM is a native 3D point-promptable part segmenter with multiple mask heads and an IoU predictor. It can be run automatically by sampling prompt points and merging redundant masks with NMS. Find3D targets open-world, language-queryable parts by auto-labeling rendered multi-view images with SAM and a VLM, projecting them back to 3D, and training a transformer to produce per-point features aligned to a CLIP-like embedding space for cosine-similarity querying. SAMPart3D and PartField both learn part-aware 3D features from multi-view SAM masks and obtain parts via feature clustering.
For interactive part segmentation, we compared our model against P3-SAM [43] and Point-SAM [86], where Point-SAM adapts the SAM prompt-and-mask paradigm to point clouds and is trained with SAM-generated pseudo masks.
Metrics. To evaluate the interactive segmentation, we sample 10 positive points for each part, then measure the average IOU between the predicted masks for all clicks of all parts and their corresponding ground truth masks. IoU@N stands for IoU score in N foreground clicks. The evaluation metric for full segmentation is the same method in previous work [43, 38], using IoU to measure the accuracy of overall mask predictions.
4.2 Main Results
4.2.1 Interactive Part-Segmentation
We evaluate interactive part segmentation on two benchmarks: PartObjaverse-Tiny [78] and PartNeXT [60]. We benchmark against two state-of-the-art native 3D methods: Point-SAM [86], which is specialized for point cloud segmentation, and P3-SAM [43]. The quantitative results are summarized in Table 1.
As shown in Table 1, SegviGen consistently outperforms all baselines by a significant margin across all interaction rounds. Notably, our method demonstrates exceptional efficiency in the few-shot interaction setting. In the most challenging $1$-click scenario (IoU@1), SegviGen achieves 42.49% on PartObjaverse-Tiny and 54.86% on PartNext, surpassing the Point-SAM by approximately 17.6% and 31.0%, respectively. This indicates that our generative framework possesses a much stronger initial understanding of 3D part structures compared to discriminative approaches, allowing it to infer complete part geometries from minimal user guidance.
Furthermore, as the number of user clicks increases from 1 to 10, SegviGen exhibits a steady and robust performance gain. On the PartNext dataset, our method reaches an IoU of 82.73% at 10 clicks, significantly higher than Point-SAM (65.04%) and P3-SAM (53.81%). This demonstrates that our model effectively incorporates user feedback to refine boundaries and resolve ambiguities.
4.2.2 Full Segmentation
We evaluate the full segmentation capability of SegviGen in two distinct settings: (1) Using purely native 3D representation. (2) Incorporating with 2D guidance. Quantitative comparisons with state-of-the-art methods, including Find3D [44], SAMPart3D [79], PartField [38], and P3-SAM [43], are presented in Table 2. Qualitative results are shown in 4.
Without 2D Guidance. In this setting, SegviGen performs segmentation solely based on the structural and appearance priors learned during pretraining, without access to any external 2D segmentation maps. The model is prompted to generate part-indicative colors directly from the latent 3D representation. As shown in Table 2, our method demonstrates superior generalization, particularly on PartNext. SegviGen achieves an IoU of 55.40%, significantly outperforming PartField (41.50%) and SAMPart3D (29.62%) While SAMPart3D performs well on the smaller PartObjaverse-Tiny dataset (59.05%), its performance collapses on PartNext. In contrast, SegviGen maintains robust performance (50.64% on PartObjaverse-Tiny).
With 2D Guidance. To further unleash the potential of SegviGen, we introduce a 2D-guided mode where the model is conditioned on a single-view 2D segmentation map (rendered via nvdiffrast or derived from a 2D segmenter). This setting combines the rich semantic cues of 2D foundation models with the geometric consistency of our 3D generative framework. Incorporating this lightweight 2D prior yields substantial performance gains. As shown in Table 2, SegviGen (w. 2D Map) achieves new state-of-the-art results on both datasets, reaching 62.98% on PartObjaverse-Tiny and 71.53% on PartNext.
Table 2: Quantitative results (IoU) for full segmentation. “w. 2D Map” denotes the setting with 2D segmentation-map guidance.
Method PartObjaverse-Tiny PartNext
Find3D [44] 15.62 19.04
SAMPart3D [79] 59.05 29.62
PartField [38] 51.72 41.50
P3-SAM [43] 45.36 31.94
SegviGen 50.64 55.40
SegviGen (w. 2D Map) 62.98 71.53
Figure 5: More qualitative segmentation results of our SegviGen.
Figure 6: Interactive 3D editing results with SegviGen and VoxHammer [27]. SegviGen provides precise part segmentation to facilitate downstream editing models. The target region to be edited is indicated by green points. It demonstrates the practical utility of SegviGen in downstream 3D editing pipelines.
4.3 Ablation Studies and Analysis
4.3.1 Point Embedding Mechanism
To investigate the optimal representation point prompt within our framework, we conducted an ablation study on the point embedding mechanism, comparing two distinct strategies:
Explicit Coordinate Encoding. In this setting, spatial coordinates are explicitly injected into the feature space. We utilize a frequency-based positional encoding scheme to map continuous 3D coordinates into high-dimensional embeddings, which will fuse with learnable semantic vectors. Consequently, the input features explicitly encapsulate both absolute spatial information and semantic category.
Label-based Semantic Embedding. In this setting, the feature vectors serve solely as semantic indicators without explicitly encoding geometric values. A shared learnable embedding vector is assigned to all foreground points. The spatial information is preserved implicitly via the coordinate indices of the SparseTensor, relying on the sparse backbone’s intrinsic ability to process spatial locality.
Table 3: Ablation study on point embedding mechanisms. We compare the performance of Explicit Coordinate Encoding and Label-based Semantic Embedding under varying numbers of clicks on PartObjaverse [78].
Method IoU@1 IoU@3 IoU@5 IoU@7 IoU@10
Explicit Coord 41.75 60.19 67.43 71.61 75.40
Label-based 42.49 61.14 67.53 71.50 75.02
Table 4: Effect of sampling steps on segmentation performance. We choose 12 steps as a practical trade-off between accuracy and efficiency.
Steps IoU@1 IoU@3 IoU@5 IoU@7 IoU@10 Time
1 42.90 59.98 65.86 69.50 72.85 0.44s
4 44.51 60.40 66.65 70.64 73.58 1.02s
8 44.21 61.14 67.64 71.14 74.49 1.81s
12 42.49 61.14 67.53 71.50 75.02 2.63s
25 43.82 61.67 68.30 71.87 74.99 5.12s
Quantitative results demonstrate a distinct trend in the performance of point embedding mechanisms. As shown in Table 3, the Label-based Semantic Embedding achieves slightly higher IoU scores when the number of clicks is limited (e.g., at 1 click). However, as the number of interactions increases, the Explicit Coordinate Encoding method consistently outperforms the Label-based approach, particularly in the later stages (e.g., 10 clicks). We attribute this phenomenon to the nature of the embeddings: while label-based embeddings are sufficient for sparse guidance, explicit positional encodings provide a finer granularity of spatial differentiation. This capability becomes increasingly critical when processing a larger set of points, allowing the model to better resolve complex geometric details from multiple user constraints.
4.3.2 Number of Denoising Steps at Inference
We analyze the impact of sampling steps on segmentation performance in Table 4. Due to the trajectory property of the flow model, we observe a great performance even with one step. Performance improves as steps increase, but gains begin to saturate over 8 steps. Although 25 steps offer marginal improvements, the inference latency nearly doubles compared to 12 steps. Thus we adopt 12 steps as the optimal balance between high-quality results and computational efficiency.
5 Conclusion
This paper introduces SegviGen, a framework that repurposes pretrained 3D generative models for 3D part segmentation. In contrast to 2D-to-3D lifting methods that often suffer from cross-view inconsistency and blurred boundaries, and native 3D discriminative approaches that require large-scale part annotations and heavy training, SegviGen transfers generative priors to deliver accurate and globally coherent segmentations with limited supervision. It reformulates segmentation as part-wise colorization, jointly reconstructing geometry and predicting part-indicative colors, and supports multiple task settings via flexible conditioning. Experiments on interactive and full segmentation benchmarks show consistent improvements over prior methods, underscoring the effectiveness and data efficiency of 3D generative priors for 3D part segmentation.
Figure 7: Interactive demo. Users specify clicks on the 3D asset to perform interactive part segmentation, and can adjust visualization settings.
Figure 8: More qualitative comparisons for full segmentation.

### 16 Apr 2026 v2

@@
 Despite encouraging progress, these native 3D approaches are fundamentally bottlenecked by the availability of large-scale, high-quality 3D part annotations, and the inconsistency of part taxonomies and granularity across datasets often introduces supervision mismatch, ultimately weakening cross-domain generalization.
+There are also related works that focus on generating compositional and complete 3D parts.
+For example, HoloPart [80] assumes segmented parts as input for part completion, Dual Volume Packing [57] directly generates complete and separable parts from a single image, and BANG [86] recovers parts by first producing an exploded mesh sequence and then applying connected-component analysis and trajectory tracking.
+In contrast, SegviGen focuses on part segmentation, which often lays the foundation for such downstream tasks.
 Figure 2: Pipeline of SegviGen.
@@
 From the results, SegviGen achieves high-accuracy full segmentation with sharp part boundaries using only 3D input.
-When 2D guidance is provided, it further allows explicit control over granularity and labels, enabling controllable, ultra-fine-grained 3D part parsing.
+With 2D guidance, it further enables explicit control over granularity and labels for ultra-fine-grained 3D part parsing.
 Implementation Details.
@@
 Baselines.
-We compared our model’s performance on full segmentation between P3-SAM [43], Find3D [44], SAMPart3D [79], Partfield [38].
+For full segmentation, we compare our method with P3-SAM [43], Find3D [44], SAMPart3D [82], and PartField [38].
 P3-SAM is a native 3D point-promptable part segmenter with multiple mask heads and an IoU predictor.
@@
 SAMPart3D and PartField both learn part-aware 3D features from multi-view SAM masks and obtain parts via feature clustering.
-For interactive part segmentation, we compared our model against P3-SAM [43] and Point-SAM [86], where Point-SAM adapts the SAM prompt-and-mask paradigm to point clouds and is trained with SAM-generated pseudo masks.
+For interactive part segmentation, we compare our model with P3-SAM [43] and Point-SAM [90], the latter of which adapts the SAM prompt-and-mask paradigm to point clouds and is trained with SAM-generated pseudo masks.
 Metrics.
@@
 4.2.2 Full Segmentation
-We evaluate the full segmentation capability of SegviGen in two distinct settings: (1) Using purely native 3D representation.
-(2) Incorporating with 2D guidance.
-Quantitative comparisons with state-of-the-art methods, including Find3D [44], SAMPart3D [79], PartField [38], and P3-SAM [43], are presented in Table 2.
-Qualitative results are shown in 4.
+We evaluate the full segmentation capability of SegviGen under two settings: (1) using only native 3D representations, and (2) incorporating 2D guidance.
+Quantitative comparisons with state-of-the-art methods are reported in Table 3, while qualitative results are shown in Fig.
+4.
 Without 2D Guidance.
@@
 2D Map) achieves new state-of-the-art results on both datasets, reaching 62.98% on PartObjaverse-Tiny and 71.53% on PartNext.
+Generalization on AI-Generated Meshes.
+Compared with manually modeled meshes, AI-generated meshes typically exhibit less well-defined part structures and weaker part separability, making part segmentation more challenging.
+To further evaluate the generalization ability of our method, we conduct experiments on synthetic test meshes generated by Hunyuan3D 2.1 [59].
+Since ground-truth annotations are unavailable for these meshes, we perform a user study on 15 models with 47 experienced 3D practitioners, comparing our method against P3-SAM [43] in terms of segmentation plausibility and boundary accuracy.
+The support rates are reported in Table 2.
+Table 2: User study on AI-generated meshes. “w. 2D Map” denotes the setting with 2D segmentation-map guidance.
+[…]
+P3-SAM [43] 15.7% 26.5%
+SegviGen 29.0% 31.4%
+SegviGen (w. 2D Map) 55.3% 42.1%
 Table 3: Quantitative results (IoU) for full segmentation. “w. 2D Map” denotes the setting with 2D segmentation-map guidance.
@@
 Experiments on interactive and full segmentation benchmarks show consistent improvements over prior methods, underscoring the effectiveness and data efficiency of 3D generative priors for 3D part segmentation.
-Figure 7: Interactive demo. Users specify clicks on the 3D asset to perform interactive part segmentation, and can adjust visualization settings.
+Figure 7: Interactive demo. Users specify clicks on the 3D asset to perform interactive part segmentation.
 Figure 8: More qualitative comparisons for full segmentation.
+6 Limitation and Future Work
+Although SegviGen achieves strong overall performance, it still faces challenges when 2D guidance is unavailable.
+This mainly arises from the semantic–instance ambiguity in part segmentation, which can make it difficult for the model to produce sufficiently clear and satisfactory part boundaries.
+For future work, we believe that SegviGen can be naturally extended to other dense 3D prediction tasks by using color as a general representation of point-wise attributes.
+For instance, color could be used to encode stress magnitude for stress map prediction, or to represent per-vertex parameters associated with different bones, such as skinning weights, for skinning-related tasks.
+This indicates that the proposed paradigm may be applicable beyond segmentation to a broader class of dense 3D prediction problems.